# Relaxation and stability analysis of a third-order multiclass traffic flow model

### v1

@@ 4.2 Relaxation to the homogeneous Aw-Rascle-Zhang model
$\partial_{t}\rho+\partial_{x}\Big{(}\rho v(\rho,z,w)\Big{)}&=0,\\ \partial_{t}z+\partial_{x}\Big{(}\frac{z^{2}}{\rho}-zq(h)\Big{)}&=0,\\ \partial_{t}w+\partial_{x}\Big{(}wv(\rho,z,w)\Big{)}&=-\frac{\textup{{W}}(\rho)-w}{\varepsilon}\end{aligned}\end{cases}$ (4.12)
@@
$\partial_{t}\rho+\partial_{x}\Big{(}\rho v\big{(}\rho,z,\textup{{W}}(\rho)\big{)}\Big{)}&=\varepsilon\partial_{x}\Big{(}\rho\mu(\rho)\partial_{x}v\big{(}\rho,z,\textup{{W}}(\rho)\big{)}\Big{)}+\mathcal{O}\big{(}\varepsilon^{2}\big{)},\\ \partial_{t}z+\partial_{x}\Big{(}\frac{z^{2}}{\rho}-zq\big{(}\textup{{H}}(\rho)\big{)}\Big{)}&=\varepsilon\partial_{x}\Big{(}z\mu(\rho)\partial_{x}v\big{(}\rho,z,\textup{{W}}(\rho)\big{)}\Big{)}+\mathcal{O}\big{(}\varepsilon^{2}\big{)}\end{aligned}\end{cases}$
@@ 4.3 Analysis of the frozen characteristics
Next, we examine the behaviour close to the frozen characteristics ${\lambda_{1}\big{(}\rho,v,\textup{{H}}(\rho)\big{)}}$. To this end, the relaxation is stated with respect to the non-conservative form ($\mathcal{S}1$), which yields the following results in terms of viscosity solutions [15, 11].
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big{(}v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v&=0,\\ \partial_{t}\Big{(}h-\textup{{h}}(\rho)\Big{)}+v\partial_{x}\Big{(}h-\textup{{h}}(\rho)\Big{)}&=-\frac{\textup{{H}}(\rho)-h}{\varepsilon}\end{aligned}\end{cases}$ (4.17)
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big{(}v-q^{\prime}\big{(}\textup{{H}}(\rho)\big{)}\textup{{h}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v&=-\varepsilon\rho\gamma(\rho)(\partial_{x}v)^{2}+\mathcal{O}\big{(}\varepsilon^{2}\big{)}\end{aligned}\end{cases}$
for $\gamma(\rho)\coloneqq q^{\prime\prime}\big{(}\textup{{H}}(\rho)\big{)}\textup{{h}}^{\prime}(\rho)\rho\big{(}\textup{{H}}^{\prime}(\rho)-\textup{{h}}^{\prime}(\rho)\big{)}.$ In terms of viscosity solutions, it is equivalent to an hyperbolic system with eigenvalues
$\lambda_{1}^{\textup{HC}}(\rho,v)$ $=\Big{(}v-q^{\prime}\big{(}\textup{{H}}(\rho)\big{)}\textup{{h}}^{\prime}(\rho)\rho\Big{)}-\varepsilon\gamma(\rho)\dot{\rho},$
$\lambda_{2}^{\textup{HC}}(\rho,v)$ $=v.$
@@
=-= - $\Big{(}\textup{{H}}^{\prime}(\rho)$ $-\textup{{h}}^{\prime}(\rho)$ $\Big{)}\rho\,\partial_{x}v$ $+\mathcal{O}(\varepsilon).$
@@
00 0 $=\partial_{t}v+\Big{(}v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v$ (4.18)
@@
00 0 $=\partial_{t}\alpha+\Big{(}v-\beta(\rho)\Big{)}\partial_{x}\alpha$ (4.19)
@@
Negleting terms of order $\mathcal{O}\big{(}\varepsilon^{2}\big{)}$, the continuity equation $\partial_{t}\rho+\partial_{x}(\rho v)=0$, equation (4.18) and (4.19) form altogether the hyperbolic quasilinear form $\partial_{t}\textup{{u}}+Q(\textup{{u}})\partial_{x}\textup{{u}}=0$ with unknowns $\textup{{u}}=(\rho,v,\alpha)^{\textup{T}}$ and matrix function
@@
$\lambda_{1}^{\textup{HC}}(\rho,v)=\big{(}v-\beta(\rho)\big{)}+\varepsilon\rho\alpha\gamma(\rho),\ \ \lambda_{2}^{\textup{HC}}(\rho,v)=v\ \ \text{and}\ \ \widetilde{\lambda_{3}^{\textup{HC}}}(\rho,v)=\big{(}v-\beta(\rho)\big{)}+\varepsilon 2\rho\alpha\gamma(\rho).$
@@
Theorem 4.3 allows to properly choose functions $\textup{{H}}(\rho)$, $\textup{{h}}(\rho)$ and $q(h)$ that account for different driver’s attitudes depending on de- and acceleration. Namely, the frozen characteristic $\lambda_{1}\big{(}\rho,v,\textup{{H}}(\rho)\big{)}$ states how fast information travels backward as a function of the density. This speed depends in the TOM model also on acceleration and deacceleration as the relation
$\lambda_{1}^{\textup{HC}}(\rho,v)=\lambda_{1}\big{(}\rho,v,\textup{{H}}(\rho)\big{)}-\varepsilon\gamma(\rho)\dot{\rho}$
@@
$\lambda_{1}^{\textup{HC}}(\rho,v)$ $\leq\lambda_{1}\big{(}\rho,v,\textup{{H}}(\rho)\big{)}$ if drivers accelerate,
$\lambda_{1}^{\textup{HC}}(\rho,v)$ $\geq\lambda_{1}\big{(}\rho,v,\textup{{H}}(\rho)\big{)}$ if drivers deaccelerate.
@@ 4.4 Relaxation to a prescribed velocity of a multiclass system
where $\widetilde{\omega}$ is a driver-dependent quantity. Such systems generate a family of velocity curves $\widetilde{\mathcal{V}}(\cdot,\widetilde{\omega})$. As explained in [19, Sec. 2.3], the ARZ model is of this form. Furthermore, the generalization of the LWR model can be explained by the choice $\widetilde{\omega}=\omega=v+P(\rho)$, where this driver-dependent quantity moves with the vehicles and influences the family of velocity curves $\widetilde{\mathcal{V}}(\rho,\omega)=\omega-P(\rho)$.
@@
[…]
@@
=-= - $\textup{{h}}^{\prime}(\rho)\Big{[}\rho\partial_{\rho}\mathcal{V}_{C}(\rho,h)\partial_{x}\rho+\rho\partial_{h}\mathcal{V}_{C}(\rho,h)\partial_{x}h\Big{]}.$
@@
$\partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big{(}v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big{)}\partial_{x}v&=\frac{\mathcal{V}_{C}(\rho,h)-v}{\varepsilon},\\ \partial_{t}w+\partial_{x}(wv)&=0\end{aligned}\end{cases}$ (4.22)
@@
Theorem 4.6 states a possible choice of the functions $q$ and h such that the novel proposed model leads to results that are justified in the recent publication [13, 14]. Namely, we have a reasonable choice if the relation $q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\approx\mathcal{V}_{C}(\rho,h)-\lambda_{1}^{C}(\rho,h)$ holds approximately, where $\mathcal{V}_{C}(\cdot,h)$ is the family of velocity curves from [13] that takes hysteresis effects into account.
@@ 5 Discussion of the Riemann problem
In this section we discuss the solution to a Riemann problem with constant left $\textup{{u}}_{\ell}$ and right $\textup{{u}}_{r}$ state. As explained in Section 3, the TOM model ($\mathcal{S}1$) – ($\mathcal{S}3$) forms a Temple class system where the fields $k=2,3$, which correspond to the double eigenvalue $\lambda_{v}(v)=v$ and which travel faster than the first wave, form contacts. Since the first field is genuinely nonlinear, it can be either a rarefaction or a shock wave, which is connected by an intermediate state $\bar{\textup{{u}}}=(\bar{\rho},\bar{v},\bar{h})^{\textup{T}}$ with the contact.
@@
The solution corresponding to a rarefaction waves travels according to an integral curve $\mathcal{I}(\cdot;\textup{{u}})$ of the first field, which satisfies the ODE $\partial_{\sigma}\mathcal{I}_{1}(\sigma;\textup{{u}})=r_{1}\big{(}\mathcal{I}(\sigma;\textup{{u}})\big{)}$ for $\mathcal{I}(0;\textup{{u}})=\textup{{u}}$ provided that it holds ${\mathcal{I}(\sigma^{*};\textup{{u}}_{\ell})=(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})^{\textup{T}}(\sigma^{*};\textup{{u}}_{\ell})=\bar{\textup{{u}}}}$ for some ${\sigma^{*}\geq 0}$. Since the intermediate states $\bar{v}$ and $\bar{h}$ are given due to the foregoing discussion, the value $\sigma^{*}$ is obtained by the equations $\mathcal{I}_{2}(\sigma^{*};\textup{{u}}_{\ell})=\bar{v}$ and $\mathcal{I}_{3}(\sigma^{*};\textup{{u}}_{\ell})=\bar{h}$. Finally, the density in the intermediate state is given by $\bar{\rho}=\mathcal{I}_{1}(\sigma^{*};\textup{{u}}_{\ell})$. According to [8, Sec. 5.1] and an entropy admissible weak solution is then given by
@@
Figure 1 – 3 show the solution to the Riemann problem with initial values $\rho_{\ell}=0.8$, $\rho_{r}=0.1$, $v_{\ell}=0.2$, $v_{r}=0.5$$h_{\ell}=\rho_{\ell}^{\nicefrac{{3}}{{2}}}$$h_{r}=\rho_{r}^{\nicefrac{{3}}{{2}}}$ and initial jump at $x_{0}=0$ for different choices of the functions $q(h)$$\textup{{h}}(\rho)$. More precisely, the panels (i) – (iii) state the density $\rho$ (blue), velocity $v$ (red) and the driver-dependent quantity $h$ (green) with respect to the left $y$-axis. The characteristic speeds $\lambda_{1}(\textup{{u}})$ are plotted with respect to the scale of the right $y$-axis. Since they are increasing, characteristics do not cross and a rarefaction wave occurs and not a shock. Panel (vi) states the difference

### v2

@@ 4.2 Relaxation to the homogeneous Aw-Rascle-Zhang model
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}\Big(\rho v(\rho,z,w)\Big)&=0,\\ \partial_{t}z+\partial_{x}\Big(\frac{z^{2}}{\rho}-zq(h)\Big)&=0,\\ \partial_{t}w+\partial_{x}\Big(wv(\rho,z,w)\Big)&=-\frac{\textup{{W}}(\rho)-w}{\varepsilon}\end{aligned}\end{cases}$ (4.12)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}\Big(\rho v\big(\rho,z,\textup{{W}}(\rho)\big)\Big)&=\varepsilon\partial_{x}\Big(\rho\mu(\rho)\partial_{x}v\big(\rho,z,\textup{{W}}(\rho)\big)\Big)+\mathcal{O}\big(\varepsilon^{2}\big),\\ \partial_{t}z+\partial_{x}\Big(\frac{z^{2}}{\rho}-zq\big(\textup{{H}}(\rho)\big)\Big)&=\varepsilon\partial_{x}\Big(z\mu(\rho)\partial_{x}v\big(\rho,z,\textup{{W}}(\rho)\big)\Big)+\mathcal{O}\big(\varepsilon^{2}\big)\end{aligned}\end{cases}$
@@ 4.3 Analysis of the frozen characteristics
Next, we examine the behaviour close to the frozen characteristics $\lambda_{1}\big(\rho,v,\textup{{H}}(\rho)\big)$}. To this end, the relaxation is stated with respect to the non-conservative form ($\mathcal{S}1$), which yields a relaxation term that is of Hamilton-Jacobi-type (HJ), where the solution is understood in terms of viscosity solutions [17, 13].
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big(v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big)\partial_{x}v&=0,\\ \partial_{t}\Big(h-\textup{{h}}(\rho)\Big)+v\partial_{x}\Big(h-\textup{{h}}(\rho)\Big)&=-\frac{\textup{{H}}(\rho)-h}{\varepsilon}\end{aligned}\end{cases}$ (4.17)
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big(v-q^{\prime}\big(\textup{{H}}(\rho)\big)\textup{{h}}^{\prime}(\rho)\rho\Big)\partial_{x}v&=-\varepsilon\rho\gamma(\rho)(\partial_{x}v)^{2}+\mathcal{O}\big(\varepsilon^{2}\big)\end{aligned}\end{cases}$
for $\gamma(\rho)\coloneqq q^{\prime\prime}\big(\textup{{H}}(\rho)\big)\textup{{h}}^{\prime}(\rho)\rho\big(\textup{{H}}^{\prime}(\rho)-\textup{{h}}^{\prime}(\rho)\big).$ In terms of viscosity solutions, it is equivalent to a hyperbolic system with eigenvalues
$\lambda_{1}^{\textup{HJ}}(\rho,v)$ $=\Big(v-q^{\prime}\big(\textup{{H}}(\rho)\big)\textup{{h}}^{\prime}(\rho)\rho\Big)-\varepsilon\gamma(\rho)\dot{\rho},$
$\lambda_{2}^{\textup{HJ}}(\rho,v)$ $=v.$
@@
$=$- $\Big(\textup{{H}}^{\prime}(\rho)$ $-\textup{{h}}^{\prime}(\rho)$ $\Big)\rho\,\partial_{x}v$ $+\mathcal{O}(\varepsilon).$
@@
$0$ $=\partial_{t}v+\Big(v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big)\partial_{x}v$ (4.18)
@@
$0$ $=\partial_{t}\alpha+\Big(v-\beta(\rho)\Big)\partial_{x}\alpha$ (4.19)
@@
Neglecting terms of order $\mathcal{O}\big(\varepsilon^{2}\big)$, the continuity equation $\partial_{t}\rho+\partial_{x}(\rho v)=0$, equation (4.18) and (4.19) form altogether the hyperbolic quasilinear form $\partial_{t}\textup{{u}}+Q(\textup{{u}})\partial_{x}\textup{{u}}=0$ with unknowns $\textup{{u}}=(\rho,v,\alpha)^{\textup{T}}$ and matrix function
@@
$\lambda_{1}^{\textup{HJ}}(\rho,v)=\big(v-\beta(\rho)\big)+\varepsilon\rho\alpha\gamma(\rho),\ \ \lambda_{2}^{\textup{HJ}}(\rho,v)=v\ \ \text{and}\ \ \widetilde{\lambda_{3}^{\textup{HJ}}}(\rho,v)=\big(v-\beta(\rho)\big)+\varepsilon 2\rho\alpha\gamma(\rho).$
@@
Theorem 4.3 allows to properly choose functions $\textup{{H}}(\rho)$, $\textup{{h}}(\rho)$ and $q(h)$ that account for different driver’s attitudes depending on de- and acceleration. Namely, the frozen characteristic $\lambda_{1}\big(\rho,v,\textup{{H}}(\rho)\big)$ states how fast information travels backward as a function of the density. This speed depends in the TOM model also on acceleration and deceleration as the relation
$\lambda_{1}^{\textup{HJ}}(\rho,v)=\lambda_{1}\big(\rho,v,\textup{{H}}(\rho)\big)-\varepsilon\gamma(\rho)\dot{\rho}$
@@
$\lambda_{1}^{\textup{HJ}}(\rho,v)$ $\leq\lambda_{1}\big(\rho,v,\textup{{H}}(\rho)\big)$ if drivers accelerate,
$\lambda_{1}^{\textup{HJ}}(\rho,v)$ $\geq\lambda_{1}\big(\rho,v,\textup{{H}}(\rho)\big)$ if drivers decelerate.
@@ 4.4 Relaxation to a prescribed velocity of a multiclass system
[…]
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}\big(\rho\mathcal{V}_{C}(\rho,h)\big)&=0,\\ \partial_{t}h+\mathcal{V}_{C}(\rho,h)\partial_{x}h&=H_{C}\big(\rho,\mathcal{V}_{C}(\rho,h)\big).\end{aligned}\end{cases}$ (4.21)
@@
$=$- $\textup{{h}}^{\prime}(\rho)\Big[\rho\partial_{\rho}\mathcal{V}_{C}(\rho,h)\partial_{x}\rho+\rho\partial_{h}\mathcal{V}_{C}(\rho,h)\partial_{x}h\Big].$
@@
$\begin{cases}\begin{aligned} \partial_{t}\rho+\partial_{x}(\rho v)&=0,\\ \partial_{t}v+\Big(v-q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\Big)\partial_{x}v&=\frac{\mathcal{V}_{C}(\rho,h)-v}{\varepsilon},\\ \partial_{t}w+\partial_{x}(wv)&=0\end{aligned}\end{cases}$ (4.22)
@@
Theorem 4.6 states a possible choice of the functions $q$ and h such that the novel proposed model leads to results that are justified in the recent publication [15, 16]. Namely, we have a reasonable choice if the relation $q^{\prime}(h)\textup{{h}}^{\prime}(\rho)\rho\approx\mathcal{V}_{C}(\rho,h)-\lambda_{1}^{C}(\rho,h)$ holds approximately, where $\mathcal{V}_{C}(\cdot,h)$ is the family of velocity curves from [15] that takes hysteresis effects into account, which can model stop-and-go waves.
@@ 5 Discussion of the Riemann problem
In this section we discuss the solution to a Riemann problem with constant left $\textup{{u}}_{\ell}$ and right $\textup{{u}}_{r}$ state. As explained in Section 3, the TOM model ($\mathcal{S}1$) – ($\mathcal{S}3$) forms a Temple class system where the fields $k=2,3$, which correspond to the double eigenvalue $\lambda_{v}(v)=v$ and which travel faster than the first wave, form contacts. Since the first field is genuinely nonlinear, it can either produce a rarefaction or a shock wave, which is connected by an intermediate state $\bar{\textup{{u}}}=(\bar{\rho},\bar{v},\bar{h})^{\textup{T}}$ with the contact.
@@
The solution corresponding to a rarefaction wave travels according to an integral curve $\mathcal{I}(\cdot;\textup{{u}})$ of the first field, which satisfies the ODE $\partial_{\sigma}\mathcal{I}_{1}(\sigma;\textup{{u}})=r_{1}\big(\mathcal{I}(\sigma;\textup{{u}})\big)$ for $\mathcal{I}(0;\textup{{u}})=\textup{{u}}$ provided that it holds $\mathcal{I}(\sigma^{*};\textup{{u}}_{\ell})=(\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3})^{\textup{T}}(\sigma^{*};\textup{{u}}_{\ell})=\bar{\textup{{u}}}$} for some $\sigma^{*}\geq 0$}. Since the intermediate states $\bar{v}$ and $\bar{h}$ are given due to the foregoing discussion, the value $\sigma^{*}$ is obtained by the equations $\mathcal{I}_{2}(\sigma^{*};\textup{{u}}_{\ell})=\bar{v}$ and $\mathcal{I}_{3}(\sigma^{*};\textup{{u}}_{\ell})=\bar{h}$. Finally, the density in the intermediate state is given by $\bar{\rho}=\mathcal{I}_{1}(\sigma^{*};\textup{{u}}_{\ell})$. According to [9, Sec. 5.1] and an entropy admissible weak solution is then given by
@@
Figures 13 show the solution to the Riemann problem with initial values $\rho_{\ell}=0.8$, $\rho_{r}=0.1$, $v_{\ell}=0.2$, $v_{r}=0.5$, $h_{\ell}=\rho_{\ell}^{\nicefrac{{3}}{{2}}}$, $h_{r}=\rho_{r}^{\nicefrac{{3}}{{2}}}$ and initial jump at $x_{0}=0$ for different choices of the functions $q(h)$, $\textup{{h}}(\rho)$. More precisely, the panels (i) – (iii) state the density $\rho$ (blue), velocity $v$ (red) and the driver-dependent quantity $h$ (green) with respect to the left $y$-axis. The characteristic speeds $\lambda_{1}(\textup{{u}})$ are plotted with respect to the scale of the right $y$-axis. Since they are increasing, characteristics do not cross and a rarefaction wave occurs and not a shock. Panel (vi) states the difference